OpenStudy (anonymous):

Newbie question. I'm doing the first questions in PSET 1, and I don't remember covering "completing the square." Its been a long time since I've done any sort of math so I'm sure this is trivial. Is this covered in the course material??

OpenStudy (anonymous):

I had the same problem, I went with Kahn's Video on it. http://www.khanacademy.org/math/algebra/quadtratics/v/completing-the-square-1?playlist=Algebra%20I%20Worked%20Examples

OpenStudy (anonymous):

Thanks for the help! It explained the "completing the square" part of the question but I still don't understand where the "translation and change of scale" come into it. I watched another one of Kahn's videos on graphing quadratics where he just threw some 'x' values and plotted the results. Is that all we're expected to do?

OpenStudy (anonymous):

Well I found someone who asked this very question and someone worked out this very problem! I found it at: http://math.stackexchange.com/questions/62075/using-translation-and-change-of-scale-to-sketch-graphs-of-these-quadratics Scroll down a bit to see the explanation. Turns out it was even easier than I expected it to be...

OpenStudy (anonymous):

Hi, I'll try to give you an answer : let \[f(x)= x ^{2}-2x\] Finding the complete square of f means you can write f as something using a square : \[f(x)=x ^{2}-2x =x ^{2}-2x+1-1\] because 1-1=0 Then you'll see a well known identity : \[x ^{3}-2x+1=(x-1)^{2}\] To conclude : \[f(x) = (x-1)^{2}-1\] Here's we can plot f from the basic function \[x ^{2}\] as follow : Translate the graph of \[x ^{2}\] by 1 to the right, and translate the resulting graph by 1 to the bottom. That's that !

OpenStudy (anonymous):

I was able to do it all the way up to your conclusion, did you shift your parabola of x^2 to the right and down by 1 because of the (x-1) and the (..-1) that follows, respectively?

OpenStudy (anonymous):

Yes exactly.